3: Determine the horizontal and vertical components of reaction at the pin A. (20pts) B A w 30°/ m m I

Use 5 m, 10 m, and 50 N/m
Answer #1

Solution - Drawing the free body diagram of the member AB we get -

W B N/m ) 60 А 30° А. 20/3 m 5 m Av m

Here value of W is -

W = \frac{1}{2}(50\: \frac{N}{m})(10 \: m) = 250 \: N

Now applying the equilibrium condition.

Taking sum of moment about point A equal to zero we get -

(R_{B})(5) - (250)(\frac{20}{3}) = 0

R_{B} = 333.333 \: N

Now taking sum of forces in horizontal direction equal to zero we get -

A_{H} - (333.333)cos(60) = 0

A_{H} = 166.667 \: (\rightarrow)

Now taking sum of forces in vertical direction equal to zero we get -

A_{V} - (333.333)sin(60) - 250 = 0

A_{V} = 38.675 \: (\uparrow)

Thus horizontal reaction at the pin A is A_{H} = 166.667 \: (\rightarrow)

And vertical reaction at the pin A is A_{V} = 38.675 \: (\uparrow)